OpenStudy (anonymous):

what is the y intercept to y=2x+3

OpenStudy (anonymous):

y-intercept is when you let x to be zero (x=0)

OpenStudy (anonymous):

like y=5x+1, let x=0 the y-intercept is: y=5(0)+1 ---> y=1